Sin^2A+ sin^2B = 1 prove that ? Please provide me solution as soon as possible 

Grade:12th pass

1 Answers

Prajwal Patankar
6 years ago
Consider a right angled Triangle ABC tight abgled at B havings sides AB, BC and AC.
Now From Pythagorus theorem, we have 
(AB)2 + (BC)= (AC)2
 
Now Divide throughout by (AC)which gives
 
(AB)2 + (BC)(AC)2
(AC)2    (AC)    (AC)2
 
(AB)2 + (BC)= 1           ... (i)
(AC)2    (AC)2
But sinA = (BC)/(AC) 
And cosA = (AB)/(AC)
 Hence, substituting values of sinA and cosA in eq(i) , we get 
 
(cosA)+ (sinA) = 1 
 
i.e. sin2A + cos2A = 1 
 
Hope it helped 👍🏻